OpenStudy (anonymous):

find the point of intersection of the two lines 2x-y=17 3x+y=13

Parth (parthkohli):

\[y = 2x - 17 \\ y = -3x + 13\]Here is what you get doing a little work.

Parth (parthkohli):

To find the intersection of two lines, just solve the system of equations.

OpenStudy (anonymous):

can u finish the problem im really confused on this id really appreciate it

geerky42 (geerky42):

Look closely. We know that \(y=2x−17\) and we also know that \(y=−3x+13\) so I can have them equal to each other, like \(\large y=2x−17 \text{ & } y=−3x+13 \Rightarrow 2x-17 = -3x + 13\) Is this clear?

geerky42 (geerky42):

\[y = 2x - 17\]\[y = -3x +13\]  \[2x - 17 = -3x + 13\]